OpenStudy (anonymous):

How can you tell when a quadratic equation has two identical, rational solutions? when the radicand is negative when b in the quadratic formula is greater than the radicand when the radicand equals zero when the radicand is not a perfect square

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (kmeis002):

If we look at the quadratic equation \[ x = \frac{-b \pm \sqrt{b-4ac}}{2a} \] Notice that the only time you will obtain different answers is when the radicand is a non zero value, since technical this gives us \[ x_{1} = \frac{-b + \sqrt{b-4ac}}{2a} \text{ or } x_{2} = \frac{-b - \sqrt{b-4ac}}{2a} \] These are two solutions that are unique for the quadratic If however \[ \sqrt{b-4ac}=0 \] Then we get: \[ x_{1} = \frac{-b}{2a} \text{ or } x_{2} = \frac{-b }{2a} \] So x_1 = x_2 when the radicand is zero

OpenStudy (anonymous):

Thank you